Suppose you just received a shipment of thirteen televisions. Three of the televisions are defective. If two televisions are randomly​ selected, compute the probability that both televisions work. What is the probability at least one of the two televisions does not​ work?
1
Expert's answer
2019-10-03T13:28:03-0400

P(both work) = ((13-3)/13) * ((13-3-1)/(13-1)) = (10*9)/(13*12) = 15/26


P(at least one does not work) = 1 - P(both work) = 1 - 15/26 = 11/26